2. (3 marks The amount invested in the term deposit was $2295. On what date was the investment made if the rate of return was 3.2% per annum? )A term deposit reached maturity on July 2,2018. The maturity value was $2345.25. (4 marks) Two payments were scheduled to be made on March 10, 2017 and December 5, 2017 of $700 and $1250 respectively. What single payment on August 15, 2017 would be equivalent to these two payments if the interest agreed upon was 6.25% per annum? 3. It 0.0 b is boter Fer deu
4. (5 marks) Seven months ago Amir agreed to pay Robert three payments of $Soo, $700 and $900 in four, twelve and eighteen months from the date of the agreement. No payments have been made and Amir and Robert have renegotiated the terms of the agreement to make one single payment six months from today. If the agreed upon rate of interest was 6% per annum. what is the size of that single payment?
Answer #1
Answer 2)
We have to provide the date of term deposit, so first
we get the total period for which interest is paid.
Thus, P (1 + r)^n = MV where r = rate of interest, n=period in years
2295 (1 + 0.032)^n = 2345.25
(1.032)^n = 2345.25/2295 = 1.021895
taking log on both side, we get
n log(1.032) = log (1.021895)
n = log (1.021895) / log(1.032)
n = 0.00940627421 / 0.0136797
n= 0.6876 year
n = 365 * 0.6876 = 251 days
Days count backward:
july 1
june 30
may 31
apr 30
march 31
feb 28
jan 31
dec 31
nov 30
oct 8
TOTAL DAYS 251
Start of deposit = 8th day of Oct, 2017 = 24th of October, 2017